Q. Consider the demand equation, Qx = 150 - PxP0 where the subscripts x and o refer to two different goods.
a. For this equation, write the expression for the point price and cross elasticities of demand as function of Px and P0.
b. what is the relationship between the price and cross elasticities ?
1
Expert's answer
2018-02-20T08:59:08-0500
a) expression for the point price for x is Ep = P0*Px/(150- P0*Px)
expression for the point price for 0 is Ep = P0*Px/(150- P0*Px)
Ecx=(P1X+P2X)/(Q1O+Q2o)*(delQ0/delPx) = -0,09(9%)
b) The higher the price you cross the elasticity is also greater.